You open your bank account on April 16 with a tax refund check of $145.00.
You pay SellUPhone $65.34 on April 18. (Check 794)
You deposit your paycheck on April 23. ($9.76/hour for 35 hours)
You buy $61.32 (4% sales tax not included) in clothes from BargainsRUs on April 24 using your debit card.
Be sure to include in your response:
the steps you followed to complete your computations
the balance as of April 24
the process for entering the transaction in the registry

Answer by josgarithmetic(39616)

Follow the money changes for the account step by step or date by date.
DATE         COMPUTATION                    BALANCE

April 16    145.00                           145.00
 
April 18    145.00-65.34=79.66                79.66

April 23     79.66+(9.76)(35)=421.26         421.26

April 24     421.26-(61.32)(1.04)=357.49     357.49